Calculate the expected percentage error

Theory says that the pendulum should follow the rule:

T=2π√(L/g)

Where T is the time period (seconds)

L is the length of the pendulum (metre)

G is the acceleration due to gravity (g = 9.81 m/s2)

Calculate the expected percentage error in the time period calculation if your measurement of length is 3% to long?
1
Expert's answer
2020-05-04T19:39:40-0400

If the percent error in the length of pendulum is "3\%" .

As,

"\frac{∆T}{T}\times100=\frac{1}{2}\frac{∆l}{l}\times100"

So, percent error in time period"=\frac{1}{2}\times3\%=1.5\%" .